OpenStudy (anonymous):

the spiffy car wash on average washes 125 cars, 75 minivans and 40 trucks weekly how many vehicles per day is this? the car wash is closed on sundays. PLZ EXPLAIN THANKS!

OpenStudy (anonymous):

125+75+40=240 240 divided by 6= 40 40 vehicles per day

OpenStudy (anonymous):

ok so take the number of cars/minivans/trucks washed and add them. So 125+75+40=240 the you divide that by the number of days they washed. that would be 6 because sunday is closed So 240/6=40 and 40 is about how many vehicles were washed that week :) please mark as helpful :)